# Average-Case Local Computation Algorithms

### v1

@@ 4 Sparse Connected Subgraphs For Sparse Erdős-Rényi
We first use Theorem 4.1 to prove the main result.
Using this distributed algorithm and the subsampling technique, we obtain the full result. See 1.9
@@
We then recall the lemma to be proved. See 4.12
@@ 4.2 Bounding Subgraph Size
By Fact 4.20 we have that with high probability every vertex has at least $\Omega(\log^{2}n)$ $2$-hop neighbors. Since a vertex being a leader implies no other member of its $2$-hop neighborhood can be a leader, (as that vertex would not have minimal index in its two-hop neighborhood) we are done. ∎
@@
where the first inequality is Fact 4.20, the second is that each administrator can be associated with at least one leader, and the third is Lemma 4.21. ∎
5 Spanners for Preferential Attachment Graphs
We next provide our LCA algorithm for sparsifying a preferential attachment graph. We first prove our structural result. As all concentration bounds hold with probability $1-n^{-c}$ for $c\geq 10$, we do not explicitly keep track of the probabilities.
5.1 Proof of Lemma 1.13
We recall the primary result which allows us to control the evolution of degrees in the graph. See 1.13 To prove the lemma, we first show a few other structural properties. We first define the degree of the PA graph after each round of sampling.
@@ 5 Spanners for Preferential Attachment Graphs
This follows from the fact that each vertex has zero degree immediately before it is added, and finishes the subsequent round with between $\mu$ and $2\mu$ edges (as in the worst case all added edges are self loops).
@@
$\Pr\left[\frac{39}{40}\sqrt{\frac{n}{t}}d_{S;t}<d_{S}<\frac{41}{40}\sqrt{\frac{n}{t}}d_{S;t}\right]\geq 1-\frac{1}{n^{10}}$
@@
where the final step follows from $d_{S,wi}\geq|S|\mu$ by Fact 5.2 and that vertices in $S$ are added in rounds below $wi$. Finally, this implies that for every edge added from $v_{i}$ in round $i$ goes to $S$ with probability at least $1-d_{S,i}/2\mu i=1-w$, and hence the probability that none of the edges are adjacent to $S$ is at most $(1-w)^{\mu}\leq n^{-100}$. ∎
@@
For convenience, let $b=\max\{c_{t},j\}$ and note that $d_{v_{i},i}\leq 2\mu$ and $d_{v_{j},\max\{c_{t},j\}}\geq d_{v_{j},j}\geq\mu$ by Fact 5.2. We have
@@
We now use Lemma 1.13 to prove Theorem 1.10. First, we establish a bound on the average query time of the algorithm, by bounding the degree of the small and large vertices.
@@
On query $(u,v)$, the algorithm checks if $v$ is the highest-degree neighbor of $u$ or vice versa, and if so keeps the edge and otherwise discards it. We first argue that this rule produces a connected subgraph of size $n-1$. By Lemma 1.13, every vertex except that of highest degree keeps an edge. Moreover, from each vertex, let $h(v)$ be the highest degree neighbor. We have that either $h(v)=v_{\max}$, the highest degree vertex, or $d_{h(v)}\geq 2\cdot d_{v}$. Thus, for every vertex $v$ the path $(v,h(v),h(h(v)),\ldots)$ has length at most $\log(n\cdot\mu)=O(\log n)$, and terminates at $v_{\max}$, with high probability. Thus, the constructed graph has diameter $O(\log n)$ as claimed.

### v2

@@ 5 Sparse Connected Subgraphs for Erdős-Rényi
We first use Theorem 5.1 to prove the main result. The algorithm of Theorem 5.1 is described formally as Algorithm 1.
Using this distributed algorithm and the subsampling technique, we obtain the full result. See 1.8
@@
We then recall the lemma to be proved. See 5.12
@@ 5.2 Bounding Subgraph Size
By Fact 5.20 we have that with high probability every vertex has at least $\Omega(\log^{2}n)$ $2$-hop neighbors. Since a vertex being a leader implies no other member of its $2$-hop neighborhood can be a leader, (as that vertex would not have minimal index in its two-hop neighborhood) we are done. ∎
@@
where the first inequality is Fact 5.20, the second is that each administrator can be associated with at least one leader, and the third is Lemma 5.21. ∎
5.3 Distributed Erdős-Rényi Sparsifier
1 Initialize $H=\{\}$
2 Round 1: for $v\in V$ do
3       $v$ sends its index to all of $v$’s neighbors
4 end for
5Round 2: for $v\in V$ do
6       if $v<u$ for all $u\in\Gamma(v)$ then
7             $v$ nominates itself as a candidate leader
8      else
9             $v$ marks itself as a non-candidate
11             $v$ adds edge $(v,\mathrm{Smallest}(v))$ to $H$
12
13       end if
14
15 end for
16Round 3: for $v\in V$, $v$ is a candidate leader do
17       if $v$ received at least one “you are not a leader” in Round 2 then
19      else
20             $v$ elects itself as a leader
21             $v$ sends “you are my administrator” to $\mathrm{Largest}(v)$
22       end if
23
24 end for
26       $v$ sets itself as an administrator
27      for $u\in\Gamma(v)$ do
28             $v$ adds $(u,v)$ to $H$
29       end for
30
31 end for
32
6 Spanners on Preferential Attachment Graphs
6.1 The Model and Related Work
We formally define the preferential attachment model:
Definition 6.1 (Preferential Attachment graph).
  • On round $1\leq i\leq n$, add a vertex $v_{i}$ into the graph.
  • Then, repeat the following process $\mu$ times:
    • The probability of the edge $(v_{i},v_{j})$ being added is $\frac{d_{v_{j}}}{\sum_{j=1}^{i}d_{v_{j}}+1}$ if $j\neq i$, and $\frac{d_{v_{j}}+1}{\sum_{j=1}^{i}d_{v_{j}}+1}$ if $j=i$.
For clarity, in this section we let $v_{i}$ for $i\in[n]$ denote the vertex that was added in round $i$. Note that the initialization of the preferential attachment graph, by the definition above, is one vertex $v_{1}$ with $\mu$ self loops.
We now describe the connections to and differences with related work. Our result is incomparable to prior work [BBC+12, FP14], which constructs an LCA [^2] Their results are written as local information algorithms (LIAs), which are sublinear algorithms with the restriction that all queries are adjacent to already explored nodes. Our results for the preferential attachment model likewise obey this restriction. The LIAs constructed in prior work find a path from any node to the root node and therefore can be interpreted as LCAs. for the root-finding problem, that of identifying a path to the first node. Their algorithm for that problem immediately implies the following:
Theorem 6.2 (Implied by [FP14]).
Let $\nu=\omega(\log n)$ be any function that dominates $\log(n)$. For every preferential attachment process with parameter $\mu$, there is an average-case LCA for $G\leftarrow G_{pa}(n,\mu)$ that w.h.p gives access to a $\nu$-spanner $H\subseteq G$. Moreover:
  • on query $(u,v)$ the LCA has time complexity $O((d_{u}+d_{v})\nu)$, which is $O(\mu\sqrt{n})$ in the worst-case and $O(\mu\operatorname{polylog}(n))$ in expectation (over all possible queries),
  • $H$ contains at most $\nu\cdot n$ edges. [^3] This bound may not be tight, but their result does not seem to give sparsity smaller than $O(n\log n)$ in any case.
In particular, note that for every degree parameter $\mu\geq\operatorname{polylog}(n)$, the algorithm obtains a $\widetilde{O}(\log n)$-spanner with $\widetilde{O}(n)$ edges. As long as the edge parameter is sufficiently large, we obtain an improved sparsity and query time bound.
Remark 6.3.
Their result is not optimized for the spanner problem, and our algorithm is essentially a simplified version of their approach tailored to this task. In particular, we may assume that the parameter $\mu$ is sufficiently large, as otherwise the graph is already sparse, whereas they solve the root-finding problem even for highly sparse graphs.
6.2 Main Structural Lemma
The proof of Theorem 1.9 relies on the following structural result about preferential attachment graphs:
Lemma 6.4.
To prove this lemma, we first show a few other structural properties. We first define the degree of the preferential attachment graph after each round of sampling.
@@ 6.2 Main Structural Lemma
This follows from the fact that each vertex has degree zero immediately before it is added, and finishes the subsequent round with between $\mu$ and $2\mu$ edges (as in the worst case all added edges are self loops).
@@
$\Pr\left[\frac{39}{40}\sqrt{\frac{n}{t}}d_{S;t}<d_{S}<\frac{41}{40}\sqrt{\frac{n}{t}}d_{S;t}\right]\geq 1-\frac{1}{n^{10}}.$
@@
where the final step follows from $d_{S,wi}\geq|S|\mu$ by Fact 6.6 and that vertices in $S$ are added in rounds below $wi$. Finally, this implies that for every edge added from $v_{i}$ in round $i$ goes to $S$ with probability at least $1-d_{S,i}/2\mu i=1-w$, and hence the probability that none of the edges are adjacent to $S$ is at most $(1-w)^{\mu}\leq n^{-100}$. ∎
@@
For convenience, let $b=\max\{c_{t},j\}$ and note that $d_{v_{i},i}\leq 2\mu$ and $d_{v_{j},\max\{c_{t},j\}}\geq d_{v_{j},j}\geq\mu$ by Fact 6.6. We have
@@
We now use Lemma 6.4 to prove Theorem 1.9. The average-case LCA is the algorithm given in Section 2.3.
First, we establish a bound on the average query time of the algorithm, by bounding the degree of the small and large vertices.
@@
On query $(u,v)$, the algorithm checks if $v$ is the highest-degree neighbor of $u$ or vice versa, and if so keeps the edge and otherwise discards it. We first argue that this rule produces a connected subgraph of size $n-1$. By Lemma 6.4, every vertex except that of highest degree keeps an edge. Moreover, from each vertex, let $h(v)$ be the highest degree neighbor. We have that either $h(v)=v_{\max}$, the highest degree vertex, or $d_{h(v)}\geq 2\cdot d_{v}$. Thus, for every vertex $v$ the path $(v,h(v),h(h(v)),\ldots)$ has length at most $\log(n\cdot\mu)$, and terminates at $v_{\max}$, with high probability. Thus, the constructed graph has diameter $O(\log n)$ as claimed.